# Online Optimization for Learning to Communicate over Time-Correlated Channels

### v1

@@ V-C Case study: $2$-ary codebook for Gaussian channels
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(C_{t})-\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\ell_{t}(C^{*})]$ (30)
$=\tilde{\mathcal{O}}\Big{(}\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sqrt{Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{(}1-Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{)}}}$
@@
(a) Baseline methods for GMD
(b) Baseline methods for LMD
(c) Channel variation of GMD
(d) Channel variation of LMD
@@
(a) Baseline methods for GMD
(b) Baseline methods for LMD
(c) Channel variation of GMD
(d) Channel variation of LMD
@@
VI Simulation results
In this section, we conduct simulation experiments to verify the empirical performance of proposed algorithms for the tasks of online decoder learning and online codebook learning, respectively. For the former, we assume that the transceiver is provided with a fixed codebook $C$ of $M$ codewords in $\mathbb{R}^{d}$, which intends to learn channel decoders $\{\mathbf{G}_{t}\}_{t\in[T]}$ for communication $T$ rounds over the time-correlated fading channel. As for the latter, the transceiver aims to select optimal codebooks $\{C_{t}\}_{t\in[T]}$ from a pre-defined super-codebook $\mathbf{C}$ with $N$ codebooks for communication $T$ rounds over the time-correlated additive noise channel. We utilize the average symbol error rate $\frac{1}{t}\sum_{\tau=1}^{t}\ell_{\tau}(D_{\tau}),t\in[T]$ defined in Eq. (5) or Eq. (9) to evaluate the performance of various methods, where $D_{\tau}$ denotes $\mathbf{G}_{\tau}$ or $C_{\tau}$. We run all experiments three times independently with different random seeds, and report the mean and standard deviation of results.
[…]
In the following, we compare the proposed algorithms with various baselines, and explore the effect of channel variation on the proposed methods for the two tasks. We set the code length $d$ and the number $M$ of codewords as $4$ below. The number $K$ of components in mixture distribution is set to $3$.
Initially, we focus on the online decoder learning task. For each round, we use a fixed codebook $C$ with the constant modulus constraint for transmission, and set the Signal-to-Noise Ratio to $24$dB. In addition, we set the degree $\rho$ of channel variation to $0.1$ for two channel distributions. For this task, we compare the proposed method with three baseline methods. The first one is to set the decoder $\mathbf{G}_{t}$ as the identity matrix $\mathbf{I}$, indicating not equalizing the channel output. The second one is to utilize the least squares estimation [32, 33] to construct $\mathbf{G}_{t}$ based on $M$ codewords $\{\mathbf{y}_{t}^{j}\}_{j\in[M]}$ received in each round. We choose Online Gradient Descent (OGD) as the third baseline method, which is widely applied in online convex optimization [20, 21]. To investigate the effect of the channel variation on the proposed method, we vary the degree $\rho$ across $\{1,3,5,10\}$ for both distributions. As illustrated in Fig. 2 (a) and (b), for two channel distributions, the proposed method shows a lower average symbol error rate compared to other baselines, and exhibits reduced variance against random channels. These results imply that the proposed method leverages distribution dependency to achieve superior performance. Moreover, from the results presented in Fig. 2 (c) and (d), the performance of optimistic OMD improves as $\rho$ decreases, suggesting that the distribution dependency assists the proposed method to learn decoders over time-correlated fading channels, matched with our theoretical findings.
(a) $M$ for GMD
(b) $M$ for LMD
Figure 4: For the online decoder learning task, we show the effect of the number $M$ of codewords on the performance of optimistic OMD.
@@ VI Simulation results
Next, we concentrate on the online codebook learning task over the time-correlated additive noise channel. The pre-constructed super-codebook $\mathbf{C}$ comprises randomly generated codebooks whose codewords are drawn from a uniform distribution in an element-wise manner. The number $N=|\mathbf{C}|$ of these codebooks is fixed as $100$. For this task, we maintain the degree $\rho$ of channel variation at $0.1$ for two distributions. Our method is compared with three baseline methods. The first baseline is to select a fixed codebook from $\mathbf{C}$ for transmission throughout the entire communication process, while the second method entails randomly selecting a codebook from $\mathbf{C}$ in each round. Additionally, we employ a classical multi-armed bandit algorithm known as Exponential-weight for Exploration and Exploitation (EXP3) [20, 28] as another baseline. To examine the impact of channel variation on our proposed method, we set the degree $\rho$ of channel variation across $\{0.5,1,3,5\}$ for conducting simulations. All experimental results are showcased in Fig. 3. In Fig. 3 (a) and (b), the proposed method demonstrates the lowest average symbol error rate over the three other baselines across the two distributions. Notably, EXP3 exhibits a similar performance with the manner of selecting a fixed codebook under Gaussian mixture distribution, and shows considerable variance under Laplace mixture distribution. It means that EXP3 overlooks utilizing the distribution dependency to boost bandit learning processes. Besides, as $\rho$ increases, signifying a more pronounced effect of channel variation, our method yields a higher average error rate, indicating that this method in fact leverages the mild environment dynamics to select the optimal codebook. These empirical observations consistently support and validate our theoretical discoveries.
We proceed to investigate the effect of the number $M$ of codewords on the performance of the proposed method for the online decoder learning task. The experimental setup is the same as the previous simulations. We vary $M$ over $\{4,8,16,32\}$, and the results for two channel distributions are shown in Fig. 4. In Fig. 4, we find that as $M$ decreases, the average error rate of the learned decoder diminishes. This observation aligns with our theoretical results for online decoder learning derived in Section IV: a smaller $M$ tightens the upper bound of expected error probability, thereby improving the performance of learned decoders.
@@ VII Conclusion
$\leq\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t})-\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t+1})+\frac{\eta_{t}}{2}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}$
@@
term (a) (35)
$\leq\sum_{t=2}^{T}\Big{(}\frac{1}{2\eta_{t}}-\frac{1}{2\eta_{t-1}}\Big{)}\|\mathbf{G}-\mathbf{G}^{\prime}_{t}\|_{F}^{2}+\frac{1}{2\eta_{1}}\|\mathbf{G}-\mathbf{G}^{\prime}_{1}\|_{F}^{2}$
@@
term (b) (36)
$=\sum_{t=2}^{T}\frac{\eta_{t}}{2}(\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2})+\|\nabla\tilde{\ell}_{1}(\mathbf{G}_{1})\|_{F}^{2}$
$\leq\sum_{t=2}^{T}\frac{\eta_{t}}{2}\Big{(}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}\Big{)}+\xi^{2}$
$=D\cdot\sum_{t=2}^{T}\frac{\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}}{2\sqrt{1+\sum_{\tau=2}^{t}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}}}+\xi^{2}$
@@
$\leq 3D+(2D+1)\xi^{2}+6Dd^{*}\sqrt{M\sum_{t=2}^{T}\sum_{j\in[M]}\big{\|}\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\big{\|}_{2}^{2}}$
@@
$=\mathcal{O}\Big{(}\sqrt{M\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{E}\big{\|}\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\big{\|}_{2}^{2}}\Big{)}.$
@@
$\leq\mathcal{O}\Big{(}\frac{1}{T}\sqrt{M\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{E}\big{\|}\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\big{\|}_{2}^{2}}\Big{)}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G})],$
where (a) holds based on the fact that for any $\mathbf{G}\in\mathcal{G}$, ~$\tilde{\ell}_{t}(\mathbf{G}_{t})-\tilde{\ell}_{t}(G)\leq\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t}),\mathbf{G}_{t}-\mathbf{G}\rangle$. This fact holds since the surrogate loss ~$\tilde{\ell}_{t}(\mathbf{G}_{t})$ is a convex function w.r.t. $\mathbf{G}_{t}$.
This completes the proof of Theorem 1
@@
$=\mathcal{O}\Big{(}M\Big{(}\sqrt{(\sigma^{\mathbf{H}}_{1:T})^{2}}+\sqrt{(\Sigma^{\mathbf{H}}_{1:T})^{2}}+\sqrt{\sigma_{W}^{2}T}\Big{)}\Big{)}.$
@@
The expected error probability $\mathbb{P}(\bar{\mathbf{G}})$ of the final output $\bar{\mathbf{G}}=\frac{1}{T}\sum_{t\in[T]}\mathbf{G}_{t}$ of Algorithm 2 satisfies
$\mathbb{P}(\bar{\mathbf{G}})$ $\leq\frac{1}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}P\Big{\{}\|\mathbf{x}^{j^{\prime}}-\bar{\mathbf{G}}\mathbf{y}^{j}\|^{2}_{2}\leq\|\mathbf{x}^{j}-\bar{\mathbf{G}}\mathbf{y}^{j}\|^{2}_{2}\Big{\}}$ (46)
$\leq\frac{1}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}\mathbb{E}\big{[}r-\|\mathbf{x}^{j^{\prime}}-\bar{\mathbf{G}}\mathbf{y}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\bar{\mathbf{G}}\mathbf{y}^{j}\|^{2}_{2}\big{]}_{+},$
where $r=2Dd^{*}L+\frac{1}{\sqrt{T}}\geq 1$ since we choose $d^{*}\geq\frac{1}{2DL}$.
Then we denote $[r-\|\mathbf{x}^{j^{\prime}}-\bar{\mathbf{G}}\mathbf{y}^{j}\|^{2}_{2}+\|\mathbf{x}^{j}-\bar{\mathbf{G}}\mathbf{y}^{j}\|^{2}_{2}]_{+}$ as ~$\tilde{\ell}^{j,j^{\prime}}(\bar{\mathbf{G}})$ and we have
$\frac{1}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}\mathbb{E}[\tilde{\ell}^{j,j^{\prime}}(\bar{\mathbf{G}})]$ $=\frac{1}{M}\sum_{j=1}^{M}\sum^{M}_{j^{\prime}\neq j}\mathbb{E}[\tilde{\ell}^{j,j^{\prime}}(\frac{1}{T}\sum_{t\in[T]}\mathbf{G}_{t})]$ (47)
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\frac{1}{MT}\sum^{M}_{j=1}\sum^{M}_{j^{\prime}\neq j}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}^{j,j^{\prime}}(\mathbf{G}_{t})]$
$=\frac{1}{MT}\sum^{M}_{j=1}\sum^{M}_{j^{\prime}\neq j}\sum_{t\in[T]}\mathbb{E}\Big{[}\tilde{\ell}^{j,j^{\prime}}(\mathbf{G}_{t})$
~$\quad-\tilde{\ell}_{t}^{j,j^{\prime}}(\mathbf{G}^{*})+\tilde{\ell}_{t}^{j,j^{\prime}}(\mathbf{G}^{*})\Big{]}$
$=\frac{\mathbb{E}[\mbox{Reg}_{T}]}{T}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})],$
where (a) holds based on Jensen’s inequality and the fact that ~$\tilde{\ell}^{j,j^{\prime}}(\bar{\mathbf{G}})$ is a convex function w.r.t $\bar{\mathbf{G}}$.
Thus we have
$\mathbb{P}(\bar{\mathbf{G}})-\frac{1}{T}\sum_{t\in T}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G}^{*})]$ $=\mathcal{O}\Big{(}M\sqrt{\frac{\sigma_{\mathbf{H}}^{2}}{T}}+M\sqrt{\frac{\sigma_{W}^{2}}{T}}\Big{)}.$ (48)
This completes the proof of Theorem 2.
-D Proof of Theorem 3
@@
$\mathbb{E}[\mbox{Reg}_{T}]=\tilde{\mathcal{O}}\Big{(}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}|\boldsymbol{\ell}_{t,i}-\boldsymbol{\ell}_{t-1,i}|}\Big{)},$ (49)
@@
Based on Lemma 3, taking the expectation on Eq. (49) over the channel noise and applying Jensen’s inequality lead to
@@
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}(C_{t})$ (53)
$\leq\frac{1}{T}\mathbb{E}[\mbox{Reg}_{T}]+\frac{1}{T}\sum_{t\in[T]}\ell_{t}(C^{*})$
@@
where $\mathbb{P}_{t}^{j}(i)=\mathbb{E}[\mathbbm{1}_{t}^{j}(i)]$.
@@
$\mathbb{E}[\mbox{Reg}_{T}]$ $=\tilde{\mathcal{O}}\Big{(}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sum_{j=1}^{2}\frac{\sigma[\mathbbm{1}_{t}^{j}(i)]}{2}}$ (54)
@@
where (a) follows from the mean value theorem of integrals and the fact that $|\exp(-x)|\leq 1,\forall x\geq 0$.
@@
$\mathbb{E}[\mbox{Reg}_{T}]$ $=\tilde{\mathcal{O}}\Big{(}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sqrt{Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{(}1-Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{)}}}$ (58)

### v3

@@ V-C Case study: $2$-ary codebook for Gaussian channels
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(C_{t})-\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\ell_{t}(C^{*})]$ (27)
$\leq\tilde{\mathcal{O}}\Big{(}\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sqrt{Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{(}1-Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{)}}}$
@@
VI Simulation Results
In this section, we conduct simulation experiments to verify the empirical performance of proposed algorithms for the tasks of online decoder learning and online codebook learning, respectively. For the former, as previously mentioned, we assume that the transmitter-receiver pair learns channel decoders $\{\mathbf{G}_{t}\}_{t\in[T]}$ via a deterministic training sequence of $M$ codewords from a fixed codebook, aimed at improving the transmission performance of data symbols transmitted over time-correlated fading channels. In the latter case, the transceiver also employs the deterministic training sequence to select optimal codebooks $\{C_{t}\}_{t\in[T]}$ from a pre-defined super-codebook $\mathbf{C}$ containing $N$ codebooks for enhancing transmission reliability of data symbols transmitted over the time-correlated additive noise channel. We utilize the average symbol error rate $\frac{1}{t}\sum_{\tau=1}^{t}\ell_{\tau}(D_{\tau}),t\in[T]$ defined in Eq. (5) or Eq. (9) to evaluate the performance of various methods, where $D_{\tau}$ denotes $\mathbf{G}_{\tau}$ or $C_{\tau}$. Building on the coherence time assumption presented in the previous sections, we can conclude that this metric reflects the performance of the learned decoder or codebook when applied to random data symbols that contain the information intended for transmission.
[…]
We first compare the proposed algorithms with various baseline methods introduced below, and set the code length $d$ as $8$ and the number $M$ of codewords as $64$ in this experiment. Then we explore the effect of channel variation on the proposed methods, and set the code length $d$ as $8$ and the number $M$ of codewords as $16$ for this experiment. The number $K$ of components in mixture distribution is set to $3$.
(a) Baselines for GMD
(b) Baselines for LMD
(c) Variation of GMD
(d) Variation of LMD
@@ VI Simulation Results
Initially, we focus on the online decoder learning task. For each round, we use a fixed codebook $C$ with the constant modulus constraint for transmission, and set the Signal-to-Noise Ratio to $24$dB. In addition, we set the degree $\rho$ of channel variation to $0.1$ for two channel distributions. For this task, we compare the proposed method with two baseline methods. The first is to utilize the least squares estimation [32] to construct $\mathbf{G}_{t}$ based on $M$ codewords $\{\mathbf{y}_{t}^{j}\}_{j\in[M]}$ received in each round. We choose Online Gradient Descent (OGD) as the second baseline method, which is widely applied in online convex optimization [18, 19]. To investigate the effect of the channel variation on the proposed method, we vary the degree $\rho$ across $\{0.1,0.3,0.5,0.7\}$ for both distributions. As illustrated in Fig. 3 (a) and (b), for two channel distributions, the proposed method shows a lower average symbol error rate compared to other baselines. Additionally, the proposed optimistic OMD method convergences faster than online gradient descent. These results imply that the proposed method leverages distribution dependency to achieve superior performance. Moreover, from the results presented in Fig. 3 (c) and (d), the performance of optimistic OMD improves as $\rho$ decreases, suggesting that the distribution dependency assists the proposed method to learn decoders over time-correlated fading channels, matched with our theoretical findings.
(a) Baselines for GMD
(b) Baselines for LMD
(c) Variation of GMD
(d) Variation of LMD
@@
Next, we concentrate on the online codebook learning task over the time-correlated additive noise channel. The pre-constructed super-codebook $\mathbf{C}$ comprises randomly generated codebooks whose codewords are drawn from a uniform distribution in an element-wise manner. The number $N=|\mathbf{C}|$ of these codebooks is fixed as $100$. For this task, we maintain the degree $\rho$ of channel variation at $0.01$ for two distributions. Our method is compared with two baseline methods. The first baseline is to randomly select a codebook from $\mathbf{C}$ in each round. Then we consider employing a classical multi-armed bandit algorithm known as Exponential-weight for Exploration and Exploitation (EXP3) [18, 29] as the second baseline. To examine the impact of channel variation on our proposed method, we set the degree $\rho$ of channel variation across $\{0.1,0.3,0.5,0.7\}$ for conducting simulations. All experimental results are showcased in Fig. 4. In Fig. 4 (a) and (b), the proposed method demonstrates the lowest average symbol error rate over the other baselines across the two distributions. Notably, EXP3 exhibits a worse performance than proposed Log-Barrier regularized optimistic OMD method. It means that EXP3 overlooks utilizing the distribution dependency to boost bandit learning processes. Besides, as $\rho$ increases, signifying a more pronounced effect of channel variation, our method yields a higher average error rate, indicating that this method in fact leverages the mild environment dynamics to select the optimal codebook. These empirical observations consistently support and validate our theoretical discoveries.
@@ VII Conclusion
$\leq\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t})-\mathcal{B}_{\psi_{t}}(\mathbf{G},\mathbf{G}^{\prime}_{t+1})+\frac{\eta_{t}}{2}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}.$
@@
term (a) $\leq\sum_{t=2}^{T}\Big{(}\frac{1}{2\eta_{t}}-\frac{1}{2\eta_{t-1}}\Big{)}\|\mathbf{G}-\mathbf{G}^{\prime}_{t}\|_{F}^{2}+\frac{1}{2\eta_{1}}\|\mathbf{G}-\mathbf{G}^{\prime}_{1}\|_{F}^{2}$ (32)
@@
term (b) $=\sum_{t=2}^{T}\frac{\eta_{t}}{2}(\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2})+\|\nabla\tilde{\ell}_{1}(\mathbf{G}_{1})\|_{F}^{2}$ (33)
$\leq D\cdot\sum_{t=2}^{T}\frac{\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}}{2\sqrt{1+\sum_{\tau=2}^{t}\|\nabla\tilde{\ell}_{t}(\mathbf{G}_{t})-\mathbf{M}_{t}\|_{F}^{2}}}+\xi^{2}$
@@
$\leq 3D+(2D+1)\xi^{2}+6Dd^{*}\sqrt{M\sum_{t=2}^{T}\sum_{j\in[M]}\big{\|}\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\big{\|}_{2}^{2}}.$
@@
$\leq\mathcal{O}\Big{(}\sqrt{M\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{E}\big{\|}\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\big{\|}_{2}^{2}}\Big{)}.$
@@
$\leq\mathcal{O}\Big{(}\sqrt{M\sum_{t\in[T]}\sum_{j\in[M]}\mathbb{E}\big{\|}\mathbf{y}_{t}^{j}-\mathbf{y}_{t-1}^{j}\big{\|}_{2}^{2}}\Big{)}+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G})]$
$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\mathcal{O}\Big{(}\frac{M}{T}\sqrt{\sum_{t\in[T]}\big{[}\gamma_{X}^{2}(\gamma_{H}^{2}-\mathbb{E}\langle\mathbf{H}_{t},\mathbf{H}_{t-1}\rangle)+\sigma_{W}^{2}\big{]}}\Big{)}$
$\quad+\frac{1}{T}\sum_{t\in[T]}\mathbb{E}[\tilde{\ell}_{t}(\mathbf{G})],$
where (a) holds based on the fact that for any $\mathbf{G}\in\mathcal{G}$, ~$\tilde{\ell}_{t}(\mathbf{G}_{t})-\tilde{\ell}_{t}(G)\leq\langle\nabla\tilde{\ell}_{t}(\mathbf{G}_{t}),\mathbf{G}_{t}-\mathbf{G}\rangle$. This fact holds since the surrogate loss ~$\tilde{\ell}_{t}(\mathbf{G}_{t})$ is a convex function w.r.t. $\mathbf{G}_{t}$. (b) holds based on the fact that $\mathbf{y}_{t}^{j}=\mathbf{H}_{t}\mathbf{x}^{j}+W_{t},\forall j\in[M].$
This completes the proof of Theorem 1.
@@
$\leq\mathcal{O}\Big{(}M\Big{(}\sqrt{(\sigma^{\mathbf{H}}_{1:T})^{2}}+\sqrt{(\Sigma^{\mathbf{H}}_{1:T})^{2}}+\sqrt{\sigma_{W}^{2}T}\Big{)}\Big{)}.$
@@
$\mathbb{E}[\mbox{Reg}_{T}]\leq\tilde{\mathcal{O}}\Big{(}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}|\boldsymbol{\ell}_{t,i}-\boldsymbol{\ell}_{t-1,i}|}\Big{)},$ (43)
@@
We first notice that the function $\sqrt{X}$ is a concave function in its domain and the Jensen’s inequality for concave function is stated as follows: if function $f(X)$ is a concave function w.r.t. the random variable $X$, then we have $\mathbb{E}[f(X)]\leq f(\mathbb{E}[X])$. Based on the above analysis and Lemma 3, taking the expectation on Eq. (43) over the channel noise and applying Jensen’s inequality lead to
@@
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(C_{t})-\frac{1}{T}\sum_{t\in[T]}\ell_{t}(C^{*})$ (47)
@@
$\quad+\frac{1}{T}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sum_{j\in[M]}\frac{\sigma[\mathbbm{1}_{t}^{j}(i)]}{M}}\Big{)},$
where $\mathbb{P}_{t}^{j}(i)=\mathbb{E}[\mathbbm{1}_{t}^{j}(i)]$.
Based on the Pinsker’s inequality [29], we can further have $|\mathbb{P}^{j}_{t}(i)-\mathbb{P}^{j}_{t-1}(i)|\leq\sqrt{\frac{1}{2}D_{kl}(\mathbb{P}^{j}_{t}(i)\|\mathbb{P}^{j}_{t-1}(i))}$ , where $D_{kl}(P\|Q):=\mathbb{E}_{P}\log\frac{dO}{dQ}$ denotes the relative entropy between probability measures $P$ and $Q$. Then we have
$\frac{1}{T}\sum_{t\in[T]}\mathbb{P}_{t}(C_{t})$ (48)
$\leq\tilde{\mathcal{O}}\Big{(}\frac{1}{T}\sqrt{N\sum_{t\in[T]}\sqrt{D_{kl}(\mathbb{P}_{t}\|\mathbb{P}_{t-1})}}$
@@
where $D_{kl}(\mathbb{P}_{t}\|\mathbb{P}_{t-1}):=\arg\max_{i,j}D_{kl}(\mathbb{P}^{j}_{t}(i)\|\mathbb{P}^{j}_{t-1}(i))$.
@@
$\mathbb{E}[\mbox{Reg}_{T}]$ $\leq\tilde{\mathcal{O}}\Big{(}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sum_{j=1}^{2}\frac{\sigma[\mathbbm{1}_{t}^{j}(i)]}{2}}$ (49)
@@
where (a) follows from the mean value theorem of integrals and $|\exp(-x)|\leq 1,\forall x\geq 0$.
@@
$\mathbb{E}[\mbox{Reg}_{T}]$ $\leq\tilde{\mathcal{O}}\Big{(}\sqrt{\sum_{i\in[N]}\sum_{t\in[T]}\sqrt{Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{(}1-Q\big{(}\frac{d_{i}}{2\sigma_{Z_{t}}}\big{)}\Big{)}}}$ (53)